A person of mass 80 kg stands at the center of a rotating merry go round platform of radius 2.8m and moment of inertia 870kg ×m^2. the platform rotates without friction with angular velocity .95 rad/s. the person walks radially to the edge of the platform.

calculate the rotational kinetic energy of the system of platform plus person before and after the persons walk
Answer #1

When the person stands at the center, moment of inertia of the system = moment of inertia of the merry go round = 870 kg.m^2 (as the moment of. inertia of the. person is zero)

Angular speed, w1 = 0.95 rad/s

Moment of inertia when the person walks to tge edge = (870+Mr^2) = (870+80*2.8^2)

= 1497.2 kgm^2

Let w2 be the angular speed of the system. By conservation of angular momentum, we have,

I1w1 = I2w2 => w2 = I1w1/I2 = (870*0.95/1497.2)

= 0.55 rad/s

Thus, rotational KE when the person stands at the center, K1 = 1/2*I1*(w1)^2 = 1/2*870*(0.95)^2

= 392.59 J

When the person walks to the edge, rotational KE becomes, K2 = 1/2*I2*(w2)^2

= 1/2*1497.2*(0.55)^2

= 226.45 J